For a right triangle that has legs a and b and hypotenuse c, find the length of the side not given here.
a = square root 5
b = square root 3

Answer by mukhopadhyay(490)

c^2=a^2+b^2 (using property)
=>c^2=5+3
=>c^2=8
=>c=sqrt(8) (negative ignored because we are looking for the length)
=c=2sqrt(2)
Answer: length of the side not given is 2sqrt(2)